OpenStudy (jessiegonzales):

ALGEBRA 2 MEDAL!!! 12. Suppose a parabola has an axis of symmetry at x=-7, a maximum height of 4 and also passes throughout the point (-6,0). Write the equation of the parabola in vertex form.

OpenStudy (jessiegonzales):

i need step by step please

OpenStudy (anonymous):

the axis of symmetry is -7 and the max height is four the vertex is (-7, 4) and the parabola opens downward vertex form y = a(x - h)^2 + k [the "a" will be negative because the parabola opens downward] vertex is (h, k) y = -a(x - (-7))^2 + 4 y = -a(x + 7)^2 + 4 solve for "a" by subbing in the given point (-6, 0) 0 = -a(-6 + 7)^2 + 4 0 = -a * 1 + 4 0 = -a + 4 -4 = -a 4 = a so your equation is y = -4(x + 7)^2 + 4